# Generalizable Metric Network for Cross-domain Person Re-identification

### v1

@@ I Introduction
As the demand for video surveillance applications grows, person re-identification (Re-ID) has become an essential technique in public security to identify individuals from videos captured by different cameras [1]. Person Re-ID has gained attention from academia and industry due to its potential in various applications [2, 3, 4, 5], where images of the same person are matched across non-overlapping camera views. However, Re-ID is a difficult task due to variations in body pose, occlusion, illumination, image resolution, background, and other factors [6, 7, 8, 9]. Person Re-ID can be considered as a special case of the image retrieval task, where the goal is to find these images from a large gallery set, which have the same identity with a given query image.
@@
Deep learning has enabled excellent performance for the person re-identification task in supervised settings [11, 12, 13, 14, 15, 16]. However, the performance drops significantly when these models are applied to new domains due to the data-distribution shift between the source domain and the unseen target domain. Collecting and labeling data is expensive and time-consuming, making existing supervised methods unsuitable for real-world applications. While some unsupervised domain adaptation methods have been developed to mitigate the labeling task [17, 18, 19, 20, 21, 22, 23], they still require data collection and model re-training for new scenarios. Therefore, the domain generalization scenario remains a challenging task in person re-identification.
@@ II-B Domain Generalization
There has been a surge of interest in developing methods to tackle the domain generalization problem in classification and semantic segmentation tasks in recent years [47, 48, 49, 50, 51, 52, 53, 54]. Many of these approaches aim to mitigate the distribution discrepancy between different domains by mapping data into a common space, following the domain adaptation paradigm [53, 55, 56, 57, 58, 59, 60]. For instance, Muandet et al.  [56] propose a kernel-based optimization algorithm to learn domain-invariant features; however, it does not ensure the consistency of conditional distribution. To overcome this limitation, Zhao et al.  [55] introduce an entropy regularization term to measure the dependence between the learned features and class labels, which promotes conditional invariance of the features and enables the classifier to perform well on features from different domains.
@@ II-C The M-Net based Person Re-ID
In the early days of deep learning, several methods employed the two-stream metric structure to obtain the similarity metric of two samples in the supervised person Re-ID task [62, 63, 64, 65]. For instance, Li et al.  [64] proposed a novel filter pairing neural network to jointly handle misalignment, photometric and geometric transforms, occlusions and background clutter. Ahmed et al.  [65] computed cross-input neighborhood differences to capture local relationships between the two input images based on mid-level features from each input image. However, these methods require a pair of images for evaluation, which resulted in feeding an image into the whole network repeatedly. In recent years, most existing person Re-ID methods directly train a robust backbone network and extract sample features for evaluation.
@@ III The proposed method
Figure 2: An illustration of the proposed Generalizable Metric Network (GMN). Here we take three source domains as an example. As seen in this figure, our method yields sample-pair features based the main (backbone) network, which are leveraged to train the metric network (M-Net). Besides, the dropout-based perturbation (DP) and the pair-identity center loss ($\mathrm{L}_{pic}$) are introduced to improve the model’s generalization and discrimination. It is worth noting that the both the backbone and M-Net are used during the test process.
@@ III-A Generalizable Metric Network
where $\odot$ is the dot-product operation. If $f^{(si)}_{p}$ and $f^{(si)}_{q}$ have the same identity, $f^{(sp)}_{k}$ is considered as a positive sample-pair feature, otherwise it is considered as a negative sample-pair feature. We use $f^{(sp)}_{k}$ as the input to train the metric network. As the input to the metric network only has two classes (positive and negative), we use a binary classification loss to train the metric network, which is given by:
@@ IV Experiments
TABLE I: Comparison with SOTA methods on Market1501 (M), MSMT17 (MS) and CUHK03 (C3). The top and the bottom are results using the training set and all images on each dataset, respectively. Note that the bold is the best result.
@@
Training Sets SNR [39] CVPR2020 8.9 8.9 6.8 19.9 34.6 62.7 16.8 30.5
[…]
MetaBIN [40] CVPR2021 28.8 28.1 17.8 40.2 57.9 80.1 34.8 49.5
META [28] ECCV2022 36.3 35.1 22.5 49.9 67.5 86.1 42.1 57.0
ACL [29] ECCV2022 41.2 41.8 20.4 45.9 74.3 89.3 45.3 59.0
GMN This paper 43.2 42.1 24.4 50.9 72.3 87.1 46.6 60.0
Full Images SNR [39] CVPR2020 17.5 17.1 7.7 22.0 52.4 77.8 25.9 39.0
[…]
MetaBIN [40] CVPR2021 43.0 43.1 18.8 41.2 67.2 84.5 43.0 56.3
META [28] ECCV2022 47.1 46.2 24.4 52.1 76.5 90.5 49.3 62.9
ACL [29] ECCV2022 49.4 50.1 21.7 47.3 76.8 90.6 49.3 62.7
GMN This paper 49.5 50.1 24.8 51.0 75.9 89.0 50.1 63.4
@@ IV-B Comparison with State-of-the-art Methods
To further confirm the superiority of the proposed method, we also performed experiments on small datasets including PRID [77], GRID [78], and VIPeR [79]. The experimental results are shown in Tab. II. As can be seen in this table, based on the averaged results of the three datasets, our method increases mAP and R1 by $1.2\%$ (72.6 vs. 71.4) and $+1.8\%$ (63.1 vs. 61.5) compared to the currently SOTA method ACL. This again validates the value of our method.
TABLE II: Comparison with SOTA methods on small datasets. All methods are trained on Market1501 (M), MSMT17 (MS), CUHK03 (C3) and CUHK-SYSU (CS).
Method PRID GRID VIPeR Average
mAP
R1
mAP
R1
mAP
R1
mAP
R1
[…]
62.2
52.3
57.4
48.6
66.3
57.0
61.9
52.6
[…]
64.3
53.1
55.0
44.4
66.2
57.5
60.6
51.7
MetaBIN [40]
70.8
61.2
57.9
50.2
64.3
55.9
61.1
55.8
META [28]
71.7
61.9
60.1
52.4
68.4
61.5
66.7
58.6
ACL [29]
73.4
63.0
65.7 55.2
75.1
66.4
71.4
61.5
GMN (Ours)
75.4 66.0
64.8
54.4
77.7 69.0 72.6 63.1
@@ IV-C Ablation Study
TABLE III: Ablation results on Market1501 (M), MSMT17 (MS) and CUHK03 (C3). “A” denotes the metric network, “B” is the dropout-based perturbation, and “C” is the pair-identity center loss. Note that the model is trained on the training set of all source domains.
@@
To validate the effectiveness of the two components of the pair-identity center loss ($L_{pic}$), namely $L_{pic+}$ and $L_{pic-}$, we conduct an experiment, and the results are shown in Tab. IV. From the table, it is evident that both $L_{pic+}$ and $L_{pic-}$ have a positive impact on improving the model’s generalization. It can be seen that the boost of $L_{pic-}$ is a little more than the boost of $L_{pic+}$. This is due to the fact that the relationship between negative sample pairs is more valuable in the process of metric. Moreover, the fused result of the two components is better than the independent result on the whole, based on the averaged result of three tasks.
TABLE IV: Ablation experiments of the pair-identity center loss on Market1501 (M), MSMT17 (MS) and CUHK03 (C3).
@@ IV-D Further Analysis
Evaluation of Different Sampling Schemes for Negative Pairs. In this section, we analyze the impact of different sampling schemes for negative pairs. Specifically, we generate negative sample pairs within each domain and across different domains, and evaluate the performance of our method on each scheme. The experimental results are reported in Tab. V. In this paper, we randomly generate negative sample pairs, which means that two samples of a negative pair could be from the same domain or different domains. As shown in the table, using the random sampling scheme leads to better performance compared to using only inter-domain or intra-domain sampling.
TABLE V: Experimental results of different sampling schemes on negative pairs on Market1501 (M), MSMT17 (MS), CUHK03 (C3) and CUHK-SYSU (CS). “Inter-domain” denotes producing negative sample-pairs across different domains. “Intra-domain” is producing negative sample-pairs within each domain. In this paper, we randomly yield the negative sample-pairs in a batch.
@@
Comparison between center (C) loss and pair-identity center (PIC) loss. We compare our PIC loss with the center loss [66] that aims to make all sample’s features with the same identity become consistent. The experimental results are given in Tab. VI. According to the comparison, we find that the proposed PIC loss can obtain better performance than the typical center loss, which thanks to the fact that our task is a metric task and the PIC loss can make the sample-pair features with the same identity consistent.
TABLE VI: Experimental results of center (C) loss and pair-identity center (PIC) loss.
@@
mAP
R1
R5
R10
mAP
R1
R5
R10
C loss [66]
41.3
40.7
63.8
73.9
71.4
87.0
94.3
96.5
PIC loss
43.2 42.1 65.4 75.0 72.3 87.1 94.5 96.7
Evaluation on different schemes on generating sample-pair features. We conduct an analysis of the impact of different schemes on generating sample-pair features, as presented in Fig. 3. Our findings indicate that the dot-product operation in Eq. 3 outperforms the absolute-value operation.
@@
Analysis of hyper-parameter sensitive. In Eq. 9, there is a hyper-parameter $\lambda$ to balance the pair-identity center loss and other losses. We conducted an experiment to analyze the effect of this hyper-parameter, as shown in Tab. VII. As observed in the table, when the hyper-parameter is set to a large or small value, the performance slightly drops compared to when $\lambda=1$. Overall, the best performance is achieved when $\lambda$ is set to 1. Therefore, in this paper, we set $\lambda$ as 1 in all experiments.
TABLE VII: Experimental results of different $\lambda$ in Eq. 9 on Market1501 (M), MSMT17 (MS) and CUHK03 (C3).
@@
TABLE VIII: Experimental results of different evaluation methods in the test stage. “Feature” denotes using features extracted by our model to evaluate performance. “M-Net” is our method uses the metric net to evaluate the performance.
@@
Analysis of the test time of different evaluation schemes. In this section, we evaluate the test time of different evaluation schemes, as shown in Tab. IX. As observed in the table, on CUHK03 and Market1501 datasets, although the Metric-Net based evaluation requires more time than feature-based evaluation, the increase in test time is relatively small. Moreover, on the larger-scale dataset, MSMT17, the Metric-Net based evaluation is slightly faster compared to the feature-based evaluation. This is because the operation of obtaining similarity through M-Net is slightly slower than the original feature-only test method, but there is no need to calculate the distance between features during the test, so the M-Net based evaluation is faster than feature-based evaluation for testing large-scale datasets.
TABLE IX: The test time of different evaluation schemes. In this table, “s” is second, and “m” is minute.
@@
Analysis of different schemes used for the dropout-based perturbation. Here, we conducted an experiment to evaluate the impact of different schemes used for the dropout-based perturbation, as listed in Tab. X. In this experiment, we applied the dropout-based perturbation at the beginning of the training stage and at the middle of the process, respectively. According to the results in the table, we observed that using the dropout-based perturbation in the middle of the training stage produced better results. In contrast, if we applied this scheme at the beginning, the perturbation could become noise to the initialized model. If we train the model after some epochs, the perturbation could be beneficial for the trained model to further enhance its generalization.
TABLE X: Experimental results of different schemes used for the dropout-based perturbation. “start” denotes that we use the dropout-based perturbation scheme at the begin of the training stage. “middle” is that we use the scheme in the middle epoch of the training process.
@@
Evaluation on different positions of the dropout-based perturbation. As is well known, ResNet consists of four blocks. In this experiment, we evaluate the impact of inserting the dropout-based perturbation at different positions within the ResNet architecture. The results are reported in Tab. XI. According to the results, when the perturbation module is inserted after Block3, the performance is the best. When it appears in the shallow layers, it may have a small impact on the Metric-Net. On the other hand, in the deep layers, it may destroy the semantic information. Therefore, inserting the drop-based perturbation module in the middle position produces satisfactory performance.
TABLE XI: Experimental results of different positions of the dropout-based perturbation. “Block1” represents that inserting the module after block1 of ResNet.
@@
Performance on source domains. Here, we also report results on source domains, as shown in Tab. XII. As seen in this table, our method performs slightly worse than the baseline on the source domains. However, for the domain generalization task, overfitting to source domains should be avoided. Therefore, our method can be considered as an effective approach to alleviate the overfitting to source domains.

### v2

@@ I Introduction
As the demand for video surveillance applications grows, person re-identification (Re-ID) has become an essential technique in public security to identify individuals from videos captured by different cameras [1, 2]. Person Re-ID has gained attention from academia and industry due to its potential in various applications [3, 4, 5, 6], where images of the same person are matched across non-overlapping camera views. However, Re-ID is a difficult task due to variations in body pose, occlusion, illumination, image resolution, background, and other factors [7, 8, 9, 10]. Person Re-ID can be considered as a special case of the image retrieval task, where the goal is to find these images from a large gallery set, which have the same identity with a given query image.
@@
Deep learning has enabled excellent performance for the person re-identification task in supervised settings [12, 13, 14, 15, 16, 17]. However, the performance drops significantly when these models are applied to new domains due to the data-distribution shift between the source domain and the unseen target domain. Collecting and labeling data is expensive and time-consuming, making existing supervised methods unsuitable for real-world applications. While some unsupervised domain adaptation methods have been developed to mitigate the labeling task [18, 19, 20, 21, 22, 23, 24, 25], they still require data collection and model re-training for new scenarios. Therefore, the domain generalization scenario remains a challenging task in person re-identification.
@@ II-A Generalizable Person Re-ID
Moreover, a novel Meta Distribution Alignment (MDA) method [49] is proposed to enable them to share similar distribution in a test-time-training fashion. Ni et al.  [50] propose a pure Transformer model (termed Part-aware Transformer) for DG-ReID by designing a proxy task to mine local visual information shared by different IDs. Jiao et al.  [51] develop a new normalization scheme called Dynamically Transformed Instance Normalization (DTIN) to alleviate the drawback of IN (i.e., the limitation of eliminating discriminative patterns). Pu et al.  [52, 53] design an Adaptive Knowledge Accumulation (AKA) framework that is endowed with two crucial abilities: knowledge representation and knowledge operation [54]. Different from these methods, our method explores the generalization capability in ReID based on the metric network.
@@ II-B Domain Generalization
There has been a surge of interest in developing methods to tackle the domain generalization problem in classification and semantic segmentation tasks in recent years [55, 56, 57, 58, 59, 60, 61, 62, 63]. Many of these approaches aim to mitigate the distribution discrepancy between different domains by mapping data into a common space, following the domain adaptation paradigm [61, 64, 65, 66, 67, 68, 69]. For instance, Muandet et al.  [65] propose a kernel-based optimization algorithm to learn domain-invariant features; however, it does not ensure the consistency of conditional distribution. To overcome this limitation, Zhao et al.  [64] introduce an entropy regularization term to measure the dependence between the learned features and class labels, which promotes conditional invariance of the features and enables the classifier to perform well on features from different domains.
@@ II-C The M-Net based Person Re-ID
In the early days of deep learning [71], several methods employed the two-stream metric structure to obtain the similarity metric of two samples in the supervised person Re-ID task [72, 73, 74, 75]. For instance, Li et al.  [74] proposed a novel filter pairing neural network to jointly handle misalignment, photometric and geometric transforms, occlusions and background clutter. Ahmed et al.  [75] computed cross-input neighborhood differences to capture local relationships between the two input images based on mid-level features from each input image. However, these methods require a pair of images for evaluation, which resulted in feeding an image into the whole network repeatedly. In recent years, most existing person Re-ID methods directly train a robust backbone network and extract sample features for evaluation.
@@ III The proposed method
Figure 2: An illustration of the proposed Generalizable Metric Network (GMN). Here we take three source domains as an example. As seen in this figure, our method yields sample-pair features based the main (backbone) network, which are leveraged to train the metric network (M-Net). Besides, the dropout-based perturbation (DP) and the pair-identity center loss ($\mathrm{L}_{pic}$) are introduced to improve the model’s generalization and discrimination. It is worth noting that both backbone and M-Net are used during the test process.
@@ III-A Generalizable Metric Network
where $\odot$ is the element-wise multiplication operation. If $f^{(si)}_{p}$ and $f^{(si)}_{q}$ have the same identity, $f^{(sp)}_{k}$ is considered as a positive sample-pair feature, otherwise it is considered as a negative sample-pair feature. We use $f^{(sp)}_{k}$ as the input to train the metric network. As the input to the metric network only has two classes (positive and negative), we use a binary classification loss to train the metric network, which is given by:
@@ III-D Explanation for GMN via Existing Theory
III-E Discussion
We will clarify the distinctions between M-Net and [82, 83] from the following four aspects. Firstly, M-Net generates sample-pair features at the feature level, which helps reduce computational costs. In contrast, prior works [82] and [83] generate sample-pair images at the image level, resulting in a disparity in the positions where sample-pair features are formed. Additionally, in terms of network architecture, both [82] and [83] are based on the framework of Siamese networks, whereas our approach deviates from this paradigm. Secondly, the approach in [83] employs dropout on the difference vector, whereas we apply dropout on the feature maps. Our objective is to enhance the diversity of features to better address domain generalization tasks. Thirdly, in comparison to [82] and [83], although M-Net utilizes the same loss function, we further introduce the pair-identity center loss to generate more discriminative features. Fourthly, while [82] and [83] have already adopted metric learning approaches to address the problem of person ReID in supervised scenarios, to the best of our knowledge, we are the first to introduce metric network for tackling DG-ReID. Our motivation is to address the challenges of generalization in this context.
Moreover, in [84] and [85], the dissimilarity representation is a vector that combines the Euclidean distances of an instance to all other instances in the set. These dissimilarity representations form the dissimilarity space of the sample set. In contrast, our sample-pair space is not related to the Euclidean distance between sample pairs. We have defined new sample-pair features that can better assist the metric network in learning domain-invariant information.
Besides, compared to these methods [86, 87, 88], they use the siamese network architecture via requiring a pair of images as input. For example, when we compute the similarity between the $i$-th image and the $j$-th image &\&& the $k$-th image, the $i$-th image is required to input the backbone two times. Different from these methods, our method uses a single network as the backbone and form the sample-pair feature in the feature space, which can effectively reduce the computation cost during training and testing. Moreover, these methods are developed to solve the supervised ReID task, while our method aims to deal with the DG-ReID task.
The dropout-based method we introduced differs from the approach in [89] in two aspects. Firstly, the dropout mechanism varies. In [89], dropout in the convolutional layer involves generating a binary mask to select corresponding channels, whereas our approach, as depicted in Fig. 2 in the paper, involves randomly masking an entire portion of a channel in the feature maps. Secondly, the purpose of dropout is distinct. In [89], dropout aims to address overfitting issues, while our objective is to enable the metric network to learn more diverse feature representations. Furthermore, the primary motivation behind introducing DP in GMN is to address the generalization challenges in the DG-ReID task more effectively.
@@ IV Experiments
TABLE I: Comparison with SOTA methods on Market1501 (M), MSMT17 (MS) and CUHK03 (C3). The top and the bottom are results using the training set and all images on each dataset, respectively. Note that the bold is the best result and † indicates evaluation results based on the code released by the authors.
@@
Training Sets SNR [41] ResNet50-IN 8.9 8.9 6.8 19.9 34.6 62.7 16.8 30.5
QAConv50 [27] ResNet50 25.4 24.8 16.4 45.3 63.1 83.7 35.0 51.3
M3[28] ResNet50-IBN 34.2 34.4 16.7 37.5 61.5 82.3 37.5 51.4
MetaBIN [42] ResNet50-BIN 28.8 28.1 17.8 40.2 57.9 80.1 34.8 49.5
META [30] ResNet50-DSBN 36.3 35.1 22.5 49.9 67.5 86.1 42.1 57.0
ACL [31] ResNet50-IBN 41.2 41.8 20.4 45.9 74.3 89.3 45.3 59.0
GMN(Ours) ResNet50-IBN 43.2 42.1 24.4 50.9 72.3 87.1 46.6 60.0
Full Images SNR [41] ResNet50-IN 17.5 17.1 7.7 22.0 52.4 77.8 25.9 39.0
QAConv50 [27] ResNet50 32.9 33.3 17.6 46.6 66.5 85.0 39.0 55.0
M3[28] ResNet50-IBN 35.7 36.5 17.4 38.6 62.4 82.7 38.5 52.6
MetaBIN [42] ResNet50-BIN 43.0 43.1 18.8 41.2 67.2 84.5 43.0 56.3
META [30] ResNet50-DSBN 47.1 46.2 24.4 52.1 76.5 90.5 49.3 62.9
ACL [31] ResNet50-IBN 49.4 50.1 21.7 47.3 76.8 90.6 49.3 62.7
IL [90] ResNet50-IBN 40.4 41.0 23.4 49.3 71.4 87.4 45.1 59.2
GMN(Ours) ResNet50-IBN 49.5 50.1 24.8 51.0 75.9 89.0 50.1 63.4
@@ IV-B Comparison with State-of-the-art Methods
We compare our method with state-of-the-art (SOTA) methods, including SNR [41], QAConv50 [27], M3[28], MetaBIN [42], META [30], IL [90] and ACL [31]. The SNR technique utilizes instance normalization to extract identity-relevant features from discarded information and reintegrates them into the network to maintain high discriminability. QAConv generates query-adaptive convolutional kernels dynamically to enable local matching. Both M3L and MetaBIN are meta-learning based methods. META employs domain-specific networks to supplement domain-specific information for each domain. ACL utilizes a shared feature space to extract both domain-invariant and domain-specific features, while also dynamically discovering relationships across different domains. The backbone of all these methods is ResNet-50 combined with instance normalization, and M3L and ACL also employ the IBN structure [94], thus it is absolutely fair in this comparison.
@@
We also conduct relevant experiments on DukeMTMC-ReID and compared them with some methods mentioned in our paper. The specific results are presented in Tab. II. Generally speaking, our method achieves better results compared to other methods. It is worth noting that META [30] uses ResNet50 with DSBN [96] (i.e., using independent normalization layer for each domain) as the backbone due to the characteristic of multiple experts. Besides, IL [90] only provides the code based on ResNet-50, thus we replace ResNet-50 with the ResNet50-IBN to run the experiment so as to achieve the fair comparison.
TABLE II: Experiments conducted on DukeMTMC-ReID (D). Market1501 (M), MSMT17 (MS) and CUHK03 (C3) are utilized as source domains during training, and DukeMTMC-ReID (D) serves as the target domain during testing. The results use the training set images on each dataset, respectively. Note that the bold is the best result and † indicates evaluation results based on the code released by the authors.
@@
mAP R1 mAP R1 mAP R1   mAP   R1
Training Sets   META [30]   ResNet50-DSBN 49.4 68.8 42.6 62.7 49.2 69.0 47.1 66.8
ACL [31] ResNet50-IBN 56.4 73.0 50.9 70.3 56.5 74.3 54.6 72.5
IL  [90] ResNet50-IBN 54.8 73.7 48.3 68.5 55.7 72.8 52.9 71.7
GMN(Ours) ResNet50-IBN 55.4 72.3 52.1 70.6 58.6 74.8 55.4 72.6
To further confirm the superiority of the proposed method, we also performed experiments on small datasets including PRID [97], GRID [98], and VIPeR [99]. The experimental results are shown in Tab. III. As can be seen in this table, based on the averaged results of the three datasets, our method increases mAP and R1 by $1.2\%$ (72.6 vs. 71.4) and $+1.8\%$ (63.1 vs. 61.5) compared to the currently SOTA method ACL. This again validates the value of our method.
TABLE III: Comparison with SOTA methods on small datasets. All methods are trained on Market1501 (M), MSMT17 (MS), CUHK03 (C3) and CUHK-SYSU (CS).
Methods PRID GRID VIPeR Average
mAP R1 mAP R1 mAP R1 mAP R1
QAConv50 [27] 62.2 52.3 57.4 48.6 66.3 57.0 61.9 52.6
M3[28] 64.3 53.1 55.0 44.4 66.2 57.5 60.6 51.7
MetaBIN [42] 70.8 61.2 57.9 50.2 64.3 55.9 61.1 55.8
META [30] 71.7 61.9 60.1 52.4 68.4 61.5 66.7 58.6
ACL [31] 73.4 63.0 65.7 55.2 75.1 66.4 71.4 61.5
GMN (Ours) 75.4 66.0 64.8 54.4 77.7 69.0 72.6 63.1
@@ IV-C Ablation Study
TABLE IV: Ablation results on Market1501 (M), MSMT17 (MS) and CUHK03 (C3). “A” denotes the metric network, “B” is the dropout-based perturbation, and “C” is the pair-identity center loss. Note that the model is trained on the training set of all source domains.
@@
To validate the effectiveness of the two components of the pair-identity center loss ($L_{pic}$), namely $L_{pic+}$ and $L_{pic-}$, we conduct an experiment, and the results are shown in Tab. V. From the table, it is evident that both $L_{pic+}$ and $L_{pic-}$ have a positive impact on improving the model’s generalization. It can be seen that the boost of $L_{pic-}$ is a little more than the boost of $L_{pic+}$. This is due to the fact that the relationship between negative sample pairs is more valuable in the process of metric. Moreover, the fused result of the two components is better than the independent result on the whole, based on the averaged result of three tasks.
TABLE V: Ablation experiments of the pair-identity center loss on Market1501 (M), MSMT17 (MS) and CUHK03 (C3).
@@ IV-D Further Analysis
Evaluation of Different Sampling Schemes for Negative Pairs. In this section, we analyze the impact of different sampling schemes for negative pairs. Specifically, we generate negative sample pairs within each domain and across different domains, and evaluate the performance of our method on each scheme. The experimental results are reported in Tab. VI. In this paper, we randomly generate negative sample pairs, which means that two samples of a negative pair could be from the same domain or different domains. As shown in the table, using the random sampling scheme leads to better performance compared to using only inter-domain or intra-domain sampling.
TABLE VI: Experimental results of different sampling schemes on negative pairs on Market1501 (M), MSMT17 (MS), CUHK03 (C3) and CUHK-SYSU (CS). “Inter-domain” denotes producing negative sample-pairs across different domains. “Intra-domain” is producing negative sample-pairs within each domain. In this paper, we randomly yield the negative sample-pairs in a batch.
@@
Comparison between center (C) loss and pair-identity center (PIC) loss. We compare our PIC loss with the center loss [76] that aims to make all sample’s features with the same identity become consistent. The experimental results are given in Tab. VII. According to the comparison, we find that the proposed PIC loss can obtain better performance than the typical center loss, which thanks to the fact that our task is a metric task and the PIC loss can make the sample-pair features with the same identity consistent.
TABLE VII: Experimental results of center (C) loss and pair-identity center (PIC) loss.
@@
mAP R1 R5 R10 mAP R1 R5 R10
C loss [76] 41.3 40.7 63.8 73.9 71.4 87.0 94.3 96.5
PIC loss 43.2 42.1 65.4 75.0 72.3 87.1 94.5 96.7
Evaluation on different schemes on generating sample-pair features. We conduct an analysis of the impact of different schemes on generating sample-pair features, as presented in Fig. 3. Our findings indicate that the dot-product operation in Eq. 3 outperforms the absolute-value operation. Moreover, we also explore various strategies, including absolute values (ABS), multiplication (MUL) and addition (ADD). We follow Protocol 2 as outlined in the paper for both training and testing. The specific experimental results are presented in Tab. VIII.
TABLE VIII: Experimental results for different strategies in creating the sample-pair space on Market1501 (M), MSMT17 (MS) and CUHK03 (C3).
[…]
  mAP R1 R5   R10   mAP R1 R5   R10  mAP R1 R5  R10 mAP R1 R5 R10
ABS 40.5 39.2 62.1 72.0 21.2 46.2 60.3 66.5 71.4 87.5 94.5 96.7 44.4 57.6 72.3 78.3
MUL 41.0 41.3 61.7 71.9 21.7 45.9 60.6 66.4 69.9 85.4 93.8 96.1 44.2 57.5 72.0 78.1
ADD 7.0 5.2 13.4 20.3 8.9 18.2 33.1 40.8 22.3 27.8 52.0 64.1 12.7 17.1 32.8 41.7
Ours 43.2 42.1 65.4 75.0 24.4 50.9 64.4 70.3 72.3 87.1 94.5 96.7 46.6 60.0 74.8 80.7
@@
Analysis of hyper-parameter sensitive. In Eq. 9, there is a hyper-parameter $\lambda$ to balance the pair-identity center loss and other losses. We conducted an experiment to analyze the effect of this hyper-parameter, as shown in Tab. IX. As observed in the table, when the hyper-parameter is set to a large or small value, the performance slightly drops compared to when $\lambda=1$. Overall, the best performance is achieved when $\lambda$ is set to 1. Therefore, in this paper, we set $\lambda$ as 1 in all experiments.
TABLE IX: Experimental results of different $\lambda$ in Eq. 9 on Market1501 (M), MSMT17 (MS) and CUHK03 (C3).
@@
TABLE X: Experimental results of different evaluation methods in the test stage. “Feature” denotes using features extracted by our model to evaluate performance. “M-Net” is our method uses the metric net to evaluate the performance.
@@
Analysis of the test time of different evaluation schemes. In this section, we evaluate the test time of different evaluation schemes, as shown in Tab. XI. As observed in the table, on CUHK03 and Market1501 datasets, although the Metric-Net based evaluation requires more time than feature-based evaluation, the increase in test time is relatively small. Moreover, on the larger-scale dataset, MSMT17, the Metric-Net based evaluation is slightly faster compared to the feature-based evaluation. This is because the operation of obtaining similarity through M-Net is slightly slower than the original feature-only test method, but there is no need to calculate the distance between features during the test, so the M-Net based evaluation is faster than feature-based evaluation for testing large-scale datasets. When analyzing the testing time under different evaluation strategies, we consistently utilized GPU (GeForce RTX 3090) for the experiments. Given the consideration of algorithmic complexity, our analysis primarily focuses on two aspects: Space Complexity and Time Complexity. Regarding space complexity, the parameter count of the backbone (ResNet-50) is 25.8M, while the parameter count of our metric network is 1.18M, accounting for only 4.5% of the backbone’s parameter count. The difference is relatively small. Regarding time complexity, our analysis focuses on forward propagation and similarity calculation. First, for forward propagation, the original backbone has a time complexity of O($f_{forward}$)), while the metric network has a time complexity of O($f_{GMN}$). As for similarity calculation, our metric network can directly compute similarity, eliminating the need for calculating a similarity matrix based on Euclidean distances between features. In contrast, the original method requires computing a similarity matrix with a complexity of O($f_{Sim}$). Therefore, the time complexity of the feature-based method is O($f_{forward}$) + O($f_{Sim}$), while our method is O($f_{forward}$) + O($f_{GMN}$). Please refer to the below formulas for a detailed estimation of the time complexity.
$O(N)\sim O(f_{forward})\sim O(NlogK),$
$O(N)\sim O(f_{forward})+O(f_{GMN})\sim O(Nlog(K+k)),$
$O(f_{Sim})\sim O(n*n),$
$O(n*n)+O(N)\sim O(f_{forward})+O(f_{Sim})$
$\sim O(n*n)+O(NlogK),$
where $K$ is the parameter count of the backbone, $k$ is the parameter count of the metric network, $N$ is the product of all dimensions of the input images, equal to the number of images multiplied by height multiplied by width multiplied by the number of channels, and $n$ is the number of images multiplied by the output dimensions of the backbone. From this, we can observe that the time complexity of the feature-based method and our method is close.
TABLE XI: The test time of different evaluation schemes. In this table, “s” is second, and “m” is minute.
@@
Analysis of different schemes used for the dropout-based perturbation. Here, we conducted an experiment to evaluate the impact of different schemes used for the dropout-based perturbation, as listed in Tab. XII. In this experiment, we applied the dropout-based perturbation at the beginning of the training stage and at the middle of the process, respectively. According to the results in the table, we observed that using the dropout-based perturbation in the middle of the training stage produced better results. In contrast, if we applied this scheme at the beginning, the perturbation could become noise to the initialized model. If we train the model after some epochs, the perturbation could be beneficial for the trained model to further enhance its generalization.
TABLE XII: Experimental results of different schemes used for the dropout-based perturbation. “start” denotes that we use the dropout-based perturbation scheme at the begin of the training stage. “middle” is that we use the scheme in the middle epoch of the training process.
@@
Evaluation on different positions of the dropout-based perturbation. As is well known, ResNet consists of four blocks. In this experiment, we evaluate the impact of inserting the dropout-based perturbation at different positions within the ResNet architecture. The results are reported in Tab. XIII. According to the results, when the perturbation module is inserted after Block3, the performance is the best. When it appears in the shallow layers, it may have a small impact on the Metric-Net. On the other hand, in the deep layers, it may destroy the semantic information. Therefore, inserting the drop-based perturbation module in the middle position produces satisfactory performance.
TABLE XIII: Experimental results of different positions of the dropout-based perturbation. “Block1” represents that inserting the module after block1 of ResNet.
@@
Performance on source domains. Here, we also report results on source domains, as shown in Tab. XV. As seen in this table, our method performs slightly worse than the baseline on the source domains. However, for the domain generalization task, overfitting to source domains should be avoided. Therefore, our method can be considered as an effective approach to alleviate the overfitting to source domains.
@@
Comparison to some big ReID models. Although big ReID models often excel at learning more intricate and abstract feature representations, crucial for handling the intricate variations in person re-identification. It is noteworthy that, from the perspective of domain generalization, intricate feature representations may not adeptly address differences in unseen domains. Our proposed method enables the network to learn the fundamental distinctions among pedestrians, thereby more effectively mitigating domain gaps. Regarding the big models you mentioned, we utilize the pre-trained models provided by LUPerson [100] and APTM [101] to extract features and evaluate their generalization performance. We conduct the following experiments in Tab. XIV. We observe that APTM offers multiple pre-trained models, and some of these models are trained on datasets that align with the same dataset used for DG-ReID, which does not meet the requirements of DG-ReID. Therefore, we opt for the pre-trained model from APTM on the PA-100K dataset [102] to serve as the testing model. Note that the results for LUPerson and APTM are directly obtained by testing the features, whereas the results for our method are derived using the Protocol 2 as outlined in our paper.
TABLE XIV: Experimental results of using pretrained model from LUPerson(LUP) [100], APTM [101] and our method on Market1501 (M), MSMT17 (MS) and CUHK03 (C3).
[…]
 mAP R1 R5  R10  mAP R1 R5  R10  mAP R1 R5  R10  mAP R1 R5  R10
LUP [100] 11.8 10.1 20.8 28.6 1.9 4.6 9.7 13.1 19.0 40.1 60.3 68.6 10.9 18.3 30.3 36.8
  APTM [101] 1.5 1.4 2.9 3.6 3.5 16.5 26.6 31.4 11.0 33.2 50.6 58.5 5.3 17.0 20.3 31.2
Ours 43.2 42.1 65.4 75.0 24.4 50.9 64.4 70.3 72.3 87.1 94.5 96.7 46.6 60.0 74.8 80.7